multiply. (y^2-7)(3y^2-8y+7)

Answer by user_dude2008(1862)

(y^2-7)(3y^2-8y+7) = y^2(3y^2-8y+7) - 7(3y^2-8y+7) = 3y^4-8y^3-14y^2+56y-49